OpenStudy (anonymous):

Evaluate the function w/o using a calculator! sec 135 Ahh please help me! I need someone to explain how to do this or show work! Thank you!

OpenStudy (anonymous):

calculate \[\cos(135)\] and then take the reciprocal

OpenStudy (anonymous):



OpenStudy (anonymous):

your angle is 135 the coordinates of that point on the unit circle are \[(-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2})\] cosine is the first coordinate

OpenStudy (anonymous):

therefore \[\cos(135)=-\frac{\sqrt{2}}{2}\] and so \[\sec(135)=-\frac{2}{\sqrt{2}}=-\sqrt{2}\]

OpenStudy (anonymous):

Thanks, can you also help me with tan(240)? I'm starting to understand. First i need to find a coterminal and then sketch the reference angle?

OpenStudy (anonymous):

yes i guess. you can also find 240 on the unit circle \[60+60+60+60=240\]

OpenStudy (anonymous):



OpenStudy (anonymous):

leftt \(-\frac{1}{2}\), down \(-\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

How do you find the square roots?